OpenStudy (anonymous):

sec^3x-cos^3x/secx-cosx=sec^2x+1+cos^2x does anyone know how to prove this?

OpenStudy (anonymous):

Put parentheses please so we know what is your exact question?

OpenStudy (anonymous):

Use the formula for \[ a^3-b^3=(a-b) \left(a^2+a b+b^2\right) \]

OpenStudy (anonymous):

Apply the above formula to get \[ \sec ^3(x)-\cos ^3(x)=(\sec (x)-\cos (x)) \left(\sec ^2(x)+\cos (x) \sec (x)+\cos^2(x)\right)\] You should be able to finish it now

OpenStudy (anonymous):

What is this product \(\cos(x) \sec(x)\) equal to? Can you show that it is equal to 1